OpenStudy (anonymous):

Chris, who is 6 feet tall, is walking away from a street light pole 30 feet high at a rate of 2 feet per second. 1. How fast is his shadow increasing in length when he is 24 feet from the pole? 2. How fast is the tip of his shadow moving when he is 24 feet from the pole? Chris, who is 6 feet tall, is walking away from a street light pole 30 feet high at a rate of 2 feet per second. 1. How fast is his shadow increasing in length when he is 24 feet from the pole? 2. How fast is the tip of his shadow moving when he is 24 feet from the pole? @Mathematics

OpenStudy (precal):

Calculus problem. Rate of change. Use \[a ^{2}+b ^{2}=c ^{2}\] Take the derivative (implicit)

OpenStudy (anonymous):

what do you think about: 24^2+30^2=L^2, L=38.42 2(x)(dx/dt)=2(L)(dL/dt), (2)(24)(2)=2(38.42)(dL/dt), dL/dt=1.25 ft/S Does this look correct

OpenStudy (precal):

\[2a(da/dt)+2b(db/dt)=2c(dc/dt)\] It has been so long since I have done these types of problems.

OpenStudy (precal):

Sorry that is all I remember for now. I hope someone can help you. I need to think on this.